Fermi National Laboratory
 

Accelerator Update for October 24, 2007 - October 25, 2007

 

Wednesday October 24

The day shift began with Operations monitoring an Antiproton Source (Pbar) stack (12.2E10) and a Recycler stash (6.8E10), and with beam to the SciBooNE / MiniBooNE experiments.

Operations put a Main Injector (MI) power supply (M:V105) into its DC mode (the supply and magnet stays on all the time and doesn’t ramp up and down) at 8:20 AM. This magnet appeared to have a temperature problem and experts wanted to see if this test would make it trip on a temperature indication.

At 8:48 AM, Operations reported a problem with a Pbar Debuncher TWT (D:V1TW04) that would not reset or turn on. An expert investigated, but found nothing wrong. 15 minutes later Operations reported the same problem. The expert investigated again and replaced a power supply.

At 9:42 AM, Operations put Pbar into standby while an expert phased the momentum cooling.

At 9:49 AM, the MI power supply interlock loop tripped off due to a volts-to-ground indication at MI-20. Operators contacted an expert.

At 10:17 AM, M:V105 tripped off on an over temperature indication for the first of two times in an hour. An expert investigated.

A Pbar expert began an Accumulator study at 11:04 AM.

At 11:11 AM, the MIPS Loop tripped off. Operations reset the system.

At 11:29 AM, the Tevatron (TeV) suffered a quench at sector A1 due to a sloppy abort, which was caused by power supply problems with a kicker (PAK2). Experts investigated and made repairs.

A MI expert reported at 12:22 PM that he needed to replace the power supply’s main contactor for M:V105; the magnet wasn’t causing the temperature problem. He had the power supply back in service about two hours later.

Operations put the TeV through a dry squeeze at 2:05 PM.

Operations established beam through SwitchYard (SY) to the Meson M01 absorber at 2:22 PM. SY experts began a study.

At 2:46 PM, the TeV ramp dumped due to a power supply failure at sector C2. Operators investigated and discovered that a technician taking a measurement had plugged into the wrong test point. Operators began recovery.

TeV experts began a study at 2:49 PM.

Operations put the TeV through a dry squeeze at 3:20 PM.

Pbar experts began a reverse proton study at 3:45 PM.

SY experts completed their study at 4:23 PM.

Recycler studies ended at 8:13 PM.

Thursday October 25

The midnight shift began TeV experts conducting studies, with Pbar conducting reverse proton studies and aperture scans, with Operations monitoring the Recycler stash, and with Sci /MiniBooNE taking beam.

Operations put the TeV through dry squeeze at 2:29 AM.

A Pelletron expert began making Electron Cooling measurements around 7 AM.

The Plans for Thursday

The current plan is for the accelerators to continue with studies and commissioning.

Machine Status

Linac: A RF station (LRF2) needs tuning

Booster: They think they’ve been having vacuum trips due to instrumentation (vacuum gauge) problems. They want to tune the injection orbit and increase intensity.

Main Injector: All their controls and configuration problems have been fixed. Their DCCT might need a new op amp. They want to try their aperture scan program today and conduct a coalescing study.

Antiproton Source: They have phased their momentum cooling. Their aperture scans have revealed some restrictions.

Recycler: They are working on aligning their electron beam.

Tevatron: They are continuing to work on tuning up the ramp.

Switchyard 120: They have a safety walkthrough today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/26/2007    email Fermilab